# Improving the solver for the Balitsky-Kovchegov evolution equation with Automatic Differentiation

### v1

@@ 3.2 Storage of the dipole amplitude
The values of the dipole amplitude are stored in an array for each value of the radial variable $r$. The array is wrapped inside a class as shown in the listing 5 in A. We provide an extended listing containing also the appropriate get and set functions. The get function always returns the result of a linear interpolation since the argument may not correspond to a node in the grid. This will often be the case in the term $S\left(\sqrt{r^{2}+r_{z}^{2}-2rr_{z}\cos\phi},\eta\right)$ in Eq. (3). In this particular case, since the argument depends itself on the variable $r$ over which derivatives are evaluated, one needs to supplement the result of the interpolation with the additional term $\partial\Big{(}\sqrt{r^{2}+r_{z}^{2}-2rr_{z}\cos\phi}\Big{)}/\partial r$ to implement the differential chain rule at the level of the array,
$\Big{(}S\big{(}f(r)\big{)}\Big{)}^{\prime}=\Big{\{}S^{\prime}\big{[}i^{+}\big{]}+(S^{\prime}\big{[}i^{+}\big{]}-S^{\prime}\big{[}i^{-}\big{]})*\big{(}f(r)-i^{+}\big{)}\Big{\}}f^{\prime}(r),$ (8)
where the integers $i^{+}$ and $i^{-}$ are the nearest larger and smaller grid nodes respectively and $S^{\prime}\big{[}i^{+}\big{]}$ and $S^{\prime}\big{[}i^{-}\big{]}$ their corresponding values of the derivative of the amplitude with respect to $r$. The derivatives $S^{\prime}\big{[}i^{+}\big{]}$ and $S^{\prime}\big{[}i^{-}\big{]}$ can be directly accessed from the array, however the term $f^{\prime}(r)$ must be inferred from the derivative of the argument. An adequate formula for the second derivative has also been implemented. In the case of the improved BK equation Eq. (4) the dipole amplitude has to be stored in a two-dimensional array for all values of rapidity $\eta$ and separation $r$. Since the rapidity shift also may depend on the variable $r$, a generalized pair of get and set functions were implemented to handle the chain rule in the case of two arguments depending on $r$.
In order to ensure differentiability of the entire algorithm we replaced all non-differentiable functions by their differentiable counterparts. In particular, instead of the $\max$ function defining the shift in rapidity of the dipole amplitudes in Eq. (4) we used the expression shown in Eq. (9),
@@ 3.3 Multi-threading
In order to decrease the run time we parallelize the calculations using multi-threading openMP library. The update of the dipole amplitude for each separation $r_{xy}$ involves the evaluation of the convolution and can be performed and stored independently. Hence, the parallelization does not involve any critical sections and can be easily implemented with an appropriate pragma.
@@ 4.1 Comparison with existing software
As a test of our software we reproduced the evaluation of the dipole amplitude $S(r,\eta)$ with the results obtained with the software used in Ref. [36]. We start with the initial condition given by the MV model [19, 20]
@@
with $e$ being the Euler constant. The initial values of the parameters are $Q_{0}^{2}=6\times 0.104\,\text{GeV}^{2}$ and $\Lambda=0.241\,\text{GeV}$ that correspond to the case of a heavy nucleus. The value $0.104\,\text{GeV}^{2}$ was obtained from the fit of the MV model to the electron-proton scattering data in [30], we multiply it by factor of 6 which represents the factor $A^{1/3}$ for a heavy nucleus with the atomic number $A$. Later, we will promote $\Lambda$ and $Q_{0}$ parameters to dual numbers and calculate the first and second derivatives of the dipole amplitude $S(r,\eta)$ with respect to these parameters.
We show the comparison in Fig.1. On the left panel we plot $S(r,\eta)$ for a wide range of $\eta$. Black lines correspond to the results from the reference code [36] and color lines from our implementation. Additionally, on the right panel we show the ratio of the dipole amplitudes to demonstrate the behavior of the results for large distances. Both codes have been run with 2000 points for the radial variable $r_{z}$ and 80 points for the $\phi$ variable. The minimal and maximal range of $r$ have been chosen accordingly in both runs. We find very good agreement between our solver and the reference code at a permille level.
Figure 1: Comparison with software from Ref. [36]. The left panel shows the dipole amplitude obtained with our software plotted in color and obtained using reference software in black. The lines overlap hiding one beneath the other. In the right panel we show their ratios for different values of $\eta$.
@@ 4.2 Comparison with divided differences
Figure 2: First and second derivative with respect to the $Q_{0}$ parameter of the dipole amplitude evolved up to $\eta=10$ as a function of $r$. Solid line correspond to the result obtained using AD, while the data points were calculated by divided differences (11,12) by varying $Q_{0}$ by about 6% by carrying two additional simulations.
@@ 4.3 Performance
In this Section we comment on the performance of our implementation. In particular, we provide timings from which the overhead generated by automatic differentiation can be estimated. We repeated our experiments for three native data types with different precision and gathered the results in Table 1. We benchmarked our implementation on a two-socket AMD EPYC 7F52 16-Core Processor with hyper-threading running with 64 openMP threads. Since the dipole amplitude rapidly vanishes with increasing $r$ the usage of the float type is not recommended as it quickly generates invalid results.
The upper part of Table 1 shows the scaling with the number of discretized points of the radial variable $r$ (we name it ”size”). The discretization of angular variable $\phi$ is fixed to 80 points for all the runs. The third column shows the time in seconds of 10 steps of the evolution (that correspond to $\eta=0.5$) when no AD is used. The scaling is quadratic (increasing the size by a factor 2 increases the time fourfold) since for each $r$ value one needs to sum contributions from the whole grid. The run in long double precision (16 bytes) is the most stable, however it takes approximately four times more time than the run in double precision (8 bytes). The fourth column shows similar timings for the simulation together with the calculation of first and second derivatives with respect to the dipole size $r$. In principle, one can obtain the $\partial S(r)/\partial r$ and $\partial^{2}S(r)/\partial r^{2}$ from the dipole amplitude itself, using divided differences approach. However, this method fails quickly for larger values of $r$ as we demonstrate later in the text. Calculating these derivatives using AD increases the time by a factor 2. This can be compared with the observation that AD in fact solves simultaneously three coupled equations: for $S(r,\eta)$, $S^{\prime}(r,\eta)$ and $S^{\prime\prime}(r,\eta)$, hence the expected increase factor would be 3. One possible source of improved efficiency is the reuse of data: see for example Listing 4 for the wrapper of the exp function where only one call to the library exp function is necessary. The fifth column contains timings when we additionally evaluate first and second derivatives with respect to two initial condition parameters including the $2\times 2$ hessian matrix. In the naive approximation of the hessian matrix by divided differences for the case of two parameters, one would need to evaluate the function 6 times: the function itself, plus and minus the increment in each parameter to get the central differences for the first derivatives and the second derivatives and one additional evaluation with both parameters incremented for the off-diagonal entry of the hessian. Counting also supplemental calculations needed for the derivatives over $r$ yields an expected factor 8 with respect to the cost without AD. Yet, we observe an only four-fold increase of time between columns 3 and 5.
@@ 5.1 Fitting experimental data
where $y=Q^{2}/sx$ is elasticity and $\sqrt{s}$ is the collision energy. $F_{L}$ can be assumed to be small and neglected in the following, therefore $\sigma_{r}=\sigma_{r}(x,Q^{2})$. $F_{2}$ is estimated for different values of Bjorken $x$ and momentum transfer $Q^{2}$. On the theoretical side $F_{2}$ can be expressed as a combination of cross sections
@@
[…]
and the sum over $f$ usually extends over three lightest quarks which we assume to be massless. $\Psi^{f}$ is the photon’s wave-function and its interaction with quark flavor $f$ can be precisely calculated,
@@
The cross section for the scattering of a pair of quark and anti-quark, $\sigma^{q\bar{q}}(\bm{\bar{r}}_{\perp},x)$ is directly related to the dipole amplitude $S(\bm{\bar{r}}_{\perp},\eta)$ discussed above via
@@
where $\alpha,\beta$ are the parameters of the initial condition and other free parameters introduced in the theoretical calculation of the reduced cross section. We would like to find their optimal values which best describe the experimental measurements. In the optimization algorithm one requires the knowledge of the gradient of $\chi^{2}$ with respect to $\alpha,\beta$ in order to advance in the parameters space. The gradients of $\chi^{2}$ can be directly estimated from the gradients of $\sigma_{r}(x,Q^{2};\alpha,\beta,\dots)$ which are given automatically by our software. In Fig. 3 we show the first and second derivatives of $\sigma^{\gamma^{*}p}_{L}(x,Q^{2})$ cross section which can be immediately translated into the gradients of $\chi^{2}$ given by Eq. (19) without the need of performing any additional simulation. We used the initial condition from Eq. (10) with two differentiable parameters: $Q_{0}$ and $\Lambda$.
Figure 3: First (left panel) and second (right panel) derivatives of the longitudinal cross section $\sigma_{L}^{\gamma^{*},p}(x,Q^{2}=10\,\text{GeV}^{2})$ given by Eq. (15) with respect to the initial condition parameters. These derivatives can be directly used to estimate the gradient of the $\chi^{2}$ function during the optimization algorithm comparing theoretical predictions to experimental measurements.
@@ 5.2 Calculating Transverse Momentum Distributions (TMD)
In order to compute the cross section for some scattering processes at high energy, it is necessary to have knowledge of transverse momentum dependent distributions (TMDs), which describe the distribution of partons within the hadron [40]. At high energy, it can be demonstrated that with a number of simplifications, the TMDs can be expressed as functions of the dipole distribution $S(r,\eta)$ and its derivatives w.r.t. $r$, see [41, 42, 43, 44, 45] (some further details are provided in B). The rapidity dependence $\eta$ of the dipole distribution is obtained by solving the BK equation. In essence, the solution of the BK equation provides access to the function $S(r,\eta)$ and its associated derivatives. However, in practice, the value of $S(r,\eta)$ is only ever known at specific values of $r$ due to the discretization of the variable $r$ on the grid. Consequently, derivatives are only approximated using a number of points on the grid. For larger values of $r$ this may give rise to large discretization artefacts, as we will show below.
@@
It is worth noticing that although $S(r,\eta)$ usually quickly decreases with $r$, the $\mathcal{K}_{\pm}(r,\eta)$ decays much slower. For example, for the Gaussian form of the dipole amplitude $S(r,\eta)=\exp(-r^{2}Q_{0}^{2})$ one gets $\mathcal{K}_{+}(r,\eta)=4r^{-2}$. As $\mathcal{K}_{+}(r,\eta)$ and $\mathcal{K}_{-}(r,\eta)$ appear in convolutions over $r$, good control of numerical precision at large $r$ is necessary. In Figure 4 we show the result for the absolute values of $\mathcal{K}_{+}(r,\eta)$ and $\mathcal{K}_{-}(r,\eta)$ calculated using our software. The left panel shows the functions for the initial condition, so only derivatives of Eq. (10) are involved. We compare the AD results with the numerical estimate of the derivatives calculated from neighboring points on a $r$ grid. Solid (dashed) lines denote results obtained using a grid with 500 points (1000 points) in the $r$ variable. It is clear that numerical derivatives (red and green curves) give unreliable results for large values of $r$. This is due to the fact that the logarithmic grid becomes sparser at large values, thus increasing the numerical error of the divided difference. Comparing result from 500 and 1000 point grids, one sees the significant difference between those two. Automatic differentiation, on the other hand, is not sensitive to mesh size because it provides analytically computed values of derivatives for each node on the grid. This can be seen from the blue and yellow lines, where plain ($500$ points) and dashed ($1000$ points) distributions converge to the same values.
@@
[…]
@@ 6 Conclusions
In this work, we have demonstrated that integrating the technique of automatic differentiation into the solver of the Balitsky-Kovechegov evolution equation can yield multiple advantages. On the one hand, the knowledge of the gradients and the hessian matrix with respect to the parameters of the initial condition can simplify and speed up the fitting procedure. This should facilitate and automate the investigation of more complicated forms of initial conditions. On the other hand, we have shown that automatic differentiation applied to the variable $r$, which is the separation between quark and anti-quark in the dipole, allows to study transverse momentum dependent parton distributions at much larger distances. Thanks to the fact that automatic differentiation is able to evaluate the derivatives analytically, the behavior of $\mathcal{K}_{-}$ and $\mathcal{K}_{+}$ at large $r$ is numerically stable. In addition, what has not been mentioned so far, but should be clear from the above discussion, is that one can use automatic differentiation on the various parameters introduced in the algorithm, such as the regularization of the dipole kernel, the value at which the current coupling is frozen, etc. By promoting these parameters to dual numbers, the simulation would automatically yield the dependence of the final results on the values of these parameters, thus helping to study the sensitivity of the results to these parameters. A similar extension may be applicable to the solution of the more general JIMWLK evolution equation. The JIMWLK equation is solved for for individual Wilson lines, so the dipole size is defined only for correlations of Wilson lines. However, the dependence of the final results on the parameters of the initial condition is of great importance and the gain in performance offered by automatic differentiation may be even greater due to the higher numerical cost of solving the JIMWLK equation.

### v2

@@ 3.2 Storage of the dipole amplitude
The values of the dipole amplitude are stored in an array for each value of the radial variable $r$. We note that appropriate get and set functions had to be implemented in order to ensure the differential chain rule. The get function always returns the result of a linear interpolation since the argument may not correspond to a node in the grid. This will often be the case in the term $S\left(\sqrt{r^{2}+r_{z}^{2}-2rr_{z}\cos\phi},\eta\right)$ in Eq. (3). In this particular case, since the argument depends itself on the variable $r$ over which derivatives are evaluated, one needs to supplement the result of the interpolation with the additional term $\partial\Big(\sqrt{r^{2}+r_{z}^{2}-2rr_{z}\cos\phi}\Big)/\partial r$ to implement the differential chain rule at the level of the array,
$\Big(S\big(f(r)\big)\Big)^{\prime}\simeq\ \Big\{S^{\prime}\big[i^{+}\big]+(S^{\prime}\big[i^{+}\big]-S^{\prime}\big[i^{-}\big])*\big(f(r)-i^{+}\big)\Big\}f^{\prime}(r),$ (7)
where the integers $i^{+}$ and $i^{-}$ are the nearest larger and smaller grid nodes respectively and $S^{\prime}\big[i^{+}\big]$ and $S^{\prime}\big[i^{-}\big]$ their corresponding values of the derivative of the amplitude with respect to $r$. The derivatives $S^{\prime}\big[i^{+}\big]$ and $S^{\prime}\big[i^{-}\big]$ can be directly accessed from the array, however, the term $f^{\prime}(r)$ must be inferred from the derivative of the argument. An adequate formula for the second derivative has also been implemented. In the case of the improved BK equation Eq. (4) the dipole amplitude has to be stored in a two-dimensional array for all values of rapidity $\eta$ and separation $r$. Since the rapidity shift also may depend on the variable $r$, a generalized pair of get and set functions were implemented to handle the chain rule in the case of two arguments depending on $r$.
In order to ensure the differentiability of the entire algorithm we replaced all non-differentiable functions by their differentiable counterparts. In particular, instead of the $\max$ function defining the shift in the rapidity of the dipole amplitudes in Eq. (4) we used the expression shown in Eq. (8),
@@ 3.3 Multi-threading
In order to decrease the run time, we parallelize the calculations using multi-threading openMP library. The update of the dipole amplitude for each separation $r_{xy}$ involves the evaluation of the convolution and can be performed and stored independently. Hence, the parallelization does not involve any critical sections and can be easily implemented with an appropriate pragma.
@@ 4.1 Comparison with existing software
As a test of our software, we reproduced the evaluation of the dipole amplitude $S(r,\eta)$ with the results obtained with the software used in Ref. [37]. We start with the initial condition given by the MV model [19, 20]
@@
with $e$ being the Euler constant. The initial values of the parameters are $Q_{0}^{2}=6\times 0.104\,\text{GeV}^{2}$ and $\Lambda=0.241\,\text{GeV}$ that correspond to the case of a heavy nucleus. The value $0.104\,\text{GeV}^{2}$ was obtained from the fit of the MV model to the electron-proton scattering data in [31], we multiply it by a factor of 6 which represents the factor $A^{1/3}$ for a heavy nucleus with the $A=216$, close to the lead mass number. In phenomenological calculations, one would modify scaling $A^{1/3}$ by multiplying it by a constant $c<1$, so that the nuclear effects are not overestimated [38]. Later, we will promote $\Lambda$ and $Q_{0}$ parameters to dual numbers and calculate the first and second derivatives of the dipole amplitude $S(r,\eta)$ with respect to these parameters.
We show the comparison in Fig.1. On the left panel, we plot $S(r,\eta)$ for a wide range of $\eta$. Black lines correspond to the results from the reference code [37] and color lines from our implementation. Additionally, on the right panel, we show the ratio of the dipole amplitudes to demonstrate the behavior of the results for large distances. Both codes have been run with 2000 points for the radial variable $r_{z}$ and 80 points for the $\phi$ variable. The minimal and maximal range of $r$ have been chosen accordingly in both runs. We find very good agreement between our solver and the reference code at a permille level.
Figure 1: Comparison with software from Ref. [37]. The left panel shows the dipole amplitude obtained with our software plotted in color and obtained using reference software in black. The lines overlap hiding one beneath the other. In the right panel, we show their ratios for different values of $\eta$.
@@ 4.2 Comparison with divided differences
Figure 2: First and second derivative with respect to the $Q_{0}$ parameter of the dipole amplitude evolved up to $\eta=10$ as a function of $r$. The solid line corresponds to the result obtained using AD, while the data points were calculated by divided differences (10,11) by varying $Q_{0}$ by about 6% by carrying two additional simulations.
@@ 4.3 Performance
In this Section, we comment on the performance of our implementation. In particular, we provide timings from which the overhead generated by automatic differentiation can be estimated. We repeated our experiments for three native data types with different precision and gathered the results in Table 1. We benchmarked our implementation on a two-socket AMD EPYC 7F52 16-Core Processor with hyper-threading running with 64 openMP threads. Since the dipole amplitude rapidly vanishes with increasing $r$ the usage of the float type is not recommended as it quickly generates invalid results.
The upper part of Table 1 shows the scaling with the number of discretized points of the radial variable $r$ (we name it ”size”). The discretization of angular variable $\phi$ is fixed to 80 points for all the runs. The third column shows the time in seconds of 10 steps of the evolution (that correspond to $\eta=0.5$) when no AD is used. The scaling is quadratic (increasing the size by a factor of 2 increases the time fourfold) since for each $r$ value one needs to sum contributions from the whole grid. The run in long double precision (16 bytes) is the most stable, however, it takes approximately four times more time than the run in double precision (8 bytes). The fourth column shows similar timings for the simulation together with the calculation of first and second derivatives with respect to the dipole size $r$. In principle, one can obtain the $\partial S(r)/\partial r$ and $\partial^{2}S(r)/\partial r^{2}$ from the dipole amplitude itself, using the divided differences approach. However, this method fails quickly for larger values of $r$ as we demonstrate later in the text. Calculating these derivatives using AD increases the time by a factor of 2. This can be compared with the observation that AD in fact solves simultaneously three coupled equations: for $S(r,\eta)$, $S^{\prime}(r,\eta)$ and $S^{\prime\prime}(r,\eta)$, hence the expected increase factor would be 3. One possible source of improved efficiency is the reuse of data: see for example Listing 3 for the wrapper of the exp function where only one call to the library exp function is necessary. The fifth column contains timings when we additionally evaluate first and second derivatives with respect to two initial condition parameters including the $2\times 2$ hessian matrix. In the naive approximation of the Hessian matrix by the divided differences for the case of two parameters, one would need to evaluate the function 6 times: the function itself, plus and minus the increment in each parameter to get the central differences for the first derivatives and the second derivatives and one additional evaluation with both parameters incremented for the off-diagonal entry of the hessian. Counting also supplemental calculations needed for the derivatives over $r$ yields an expected factor 8 with respect to the cost without AD. Yet, we observe only a four-fold increase in time between columns 3 and 5.
@@ 5.1 Fitting experimental data
where $y=Q^{2}/sx$ is elasticity and $\sqrt{s}$ is the collision energy. $F_{L}$ can be assumed to be small and neglected in the following, therefore $\sigma_{r}=\sigma_{r}(x,Q^{2})$. $F_{2}$ is estimated for different values of Bjorken $x$ and momentum transfer $Q^{2}$. On the theoretical side $F_{2}$ can be expressed as a combination of cross-sections
@@
$\sigma^{\gamma^{*}p}_{T,L}(x,Q^{2})=\sum_{f}\int d^{2}\bm{\bar{r}}_{\perp}\int_{0}^{1}dz\,|\Psi^{f}_{T,L}(\bm{\bar{r}}_{\perp},z,Q^{2})|^{2}\sigma^{q\bar{q}}(\bm{\bar{r}}_{\perp},x),$ (14)
and the sum over $f$ usually extends over three lightest quarks which we assume to be massless. $\Psi^{f}$ is the photon’s wave-function and its interaction with quark flavor $f$ can be calculated in perturbation theory. At leading leading order in the QED coupling $\alpha_{em}$ (and zeroth order in $\alpha_{s}$), one has
@@
The cross-section for the scattering of a pair of quark and anti-quark, $\sigma^{q\bar{q}}(\bm{\bar{r}}_{\perp},x)$ is directly related to the dipole amplitude $S(\bm{\bar{r}}_{\perp},\eta)$ discussed above via
@@
where $\alpha,\beta$ are the parameters of the initial condition and other free parameters introduced in the theoretical calculation of the reduced cross-section. We would like to find the optimal values that best describe the experimental measurements. In the optimization algorithm one requires the knowledge of the gradient of $\chi^{2}$ with respect to $\alpha,\beta$ in order to advance in the parameters space. The gradients of $\chi^{2}$ can be directly estimated from the gradients of $\sigma_{r}(x,Q^{2};\alpha,\beta,\dots)$ which are given automatically by our software. In Fig. 3 we show the first and second derivatives of $\sigma^{\gamma^{*}p}_{L}(x,Q^{2})$ cross-section which can be immediately translated into the gradients of $\chi^{2}$ given by Eq. (18) without the need to perform any additional simulation. We used the initial condition from Eq. (9) with two differentiable parameters: $Q_{0}$ and $\Lambda$.
Figure 3: First (left panel) and second (right panel) derivatives of the longitudinal cross-section $\sigma_{L}^{\gamma^{*},p}(x,Q^{2}=10\,\text{GeV}^{2})$ given by Eq. (14) with respect to the initial condition parameters. These derivatives can be directly used to estimate the gradient of the $\chi^{2}$ function during the optimization algorithm comparing theoretical predictions to the experimental measurements.
@@ 5.2 Calculating Transverse Momentum Distributions (TMD)
In order to compute the cross section for some scattering processes at high energy, it is necessary to have the knowledge of transverse momentum dependent distributions (TMDs), which describe the distribution of partons within the hadron [42]. At high energy, it can be demonstrated that with a number of simplifications, the TMDs can be expressed as functions of the dipole distribution $S(r,\eta)$ and its derivatives w.r.t. $r$, see [43, 44, 45, 46, 47] (some further details are provided in B). The rapidity dependence $\eta$ of the dipole distribution is obtained by solving the BK equation. In essence, the solution of the BK equation provides access to the function $S(r,\eta)$ and its associated derivatives. However, in practice, the value of $S(r,\eta)$ is only ever known at specific values of $r$ due to the discretization of the variable $r$ on the grid. Consequently, derivatives are only approximated using a number of points on the grid. For larger values of $r$ this may give rise to large discretization artefacts, as we will show below.
@@
It is worth noticing that although $S(r,\eta)$ usually quickly decreases with $r$, the $\mathcal{K}_{\pm}(r,\eta)$ decays much slower. For example, for the Gaussian form of the dipole amplitude $S(r,\eta)=\exp(-r^{2}Q_{0}^{2})$ one gets $\mathcal{K}_{+}(r,\eta)=4r^{-2}$. As $\mathcal{K}_{+}(r,\eta)$ and $\mathcal{K}_{-}(r,\eta)$ appear in convolutions over $r$, good control of numerical precision at large $r$ is necessary. In Figure 4 we show the result for the absolute values of $\mathcal{K}_{+}(r,\eta)$ and $\mathcal{K}_{-}(r,\eta)$ calculated using our software. The left panel shows the functions for the initial condition, so only derivatives of Eq. (9) are involved. We compare the AD results with the numerical estimate of the derivatives calculated from neighboring points on a $r$ grid. Solid (dashed) lines denote results obtained using a grid with 500 points (1000 points) in the $r$ variable. It is clear that numerical derivatives (red and green curves) give unreliable results for large values of $r$. This is due to the fact that the logarithmic grid becomes sparser at large values, thus increasing the numerical error of the divided difference. Comparing results from 500 and 1000 point grids, one sees the significant difference between those two. Automatic differentiation, on the other hand, is not sensitive to mesh size because it provides analytically computed values of derivatives for each node on the grid. This can be seen from the blue and yellow lines, where plain ($500$ points) and dashed ($1000$ points) distributions converge to the same values.
@@
Figure 4: Left panel: functions $|\mathcal{K}_{+}(r_{\perp})|$ and $|\mathcal{K}_{-}(r_{\perp})|$ at initial condition $\eta=0$ for two grid sizes, N=500, 1000N=500,\ 1000 and two methods of evaluation: AD and divided difference on the grid points. Right panel: functions $|\mathcal{K}_{\pm}(r_{\perp})|$ after evolution to $\eta=4$ for four grid sizes: N=500, 1000, 2000, 4000N=500,\ 1000,\ 2000,\ 4000.
@@ 6 Conclusions
In this work, we have demonstrated that integrating the technique of automatic differentiation into the solver of the Balitsky-Kovchegov evolution equation can yield multiple advantages. On the one hand, the knowledge of the gradients and the hessian matrix with respect to the parameters of the initial condition can simplify and speed up the fitting procedure. This should facilitate and automate the investigation of more complicated forms of initial conditions. On the other hand, we have shown that automatic differentiation applied to the variable $r$, which is the separation between quark and anti-quark in the dipole, allows to study transverse momentum dependent parton distributions at much larger distances. Thanks to the fact that automatic differentiation is able to evaluate the derivatives analytically, the behavior of $\mathcal{K}_{-}$ and $\mathcal{K}_{+}$ at large $r$ is numerically stable. In addition, what has not been mentioned so far, but should be clear from the above discussion, is that one can use automatic differentiation on the various parameters introduced in the algorithm, such as the regularization of the dipole kernel, the value at which the current coupling is frozen, etc. By promoting these parameters to dual numbers, the simulation would automatically yield the dependence of the final results on the values of these parameters, thus helping to study the sensitivity of the results to these parameters. A similar extension may be applicable to the solution of the more general JIMWLK evolution equation. The JIMWLK equation is solved for for individual Wilson lines, so the dipole size is defined only for correlations of Wilson lines. However, the dependence of the final results on the parameters of the initial condition is of great importance and the gain in performance offered by automatic differentiation may be even greater due to the higher numerical cost of solving the JIMWLK equation.